OpenStudy (anonymous):

Factor each polynomial. 1. 4y2 + 8 2. y2 - 4 3. x2 - 14x - 15 4. 4x2 - 3x + 20x - 15. 5. x2y3 + xy 6. 3x3 + 6x2 – 24x.

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

which one would you like to factor?

OpenStudy (anonymous):

Hey girl!

OpenStudy (misty1212):

hey \(\color\magenta\heartsuit\)\(\color\magenta\heartsuit\)

OpenStudy (misty1212):

want to do \(4y^2 + 8\) first ?

OpenStudy (anonymous):

yess

OpenStudy (misty1212):

k you see a common factor in \(4y^2\) and \(8\) ? it is just a number common to both

OpenStudy (anonymous):

they both dont have a same number? lol sorry im confused

OpenStudy (misty1212):

i know they don't' one has a 4 the other has an 8 you know a way to factor 8?

OpenStudy (misty1212):

guess not they have a common factor of 4 since \(8=4\times 2\)so you can "factor it out" and you get \[4y^2+8=4(y^2+2)\]

OpenStudy (anonymous):

im sorry i was on other page of school lol and yes 8 okay cool

OpenStudy (jhannybean):

That's quite a lot of questions. Lol.

OpenStudy (misty1212):

HI @Jhannybean please step in and help i have to get up early !!

OpenStudy (jhannybean):

Me too!! What class do you have tomorrow?

OpenStudy (anonymous):

i know:( lol

OpenStudy (misty1212):

no math just stuff you?

OpenStudy (jhannybean):

Chem, haha. Oh boy :)

OpenStudy (jhannybean):

But yeah... are you done with #1 @brittney_IDKM ?

OpenStudy (misty1212):

good luck good night too `

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Alright, #2.

OpenStudy (anonymous):

is number 2 (x+2 )(y-2)?

OpenStudy (jhannybean):

\[y^2-4 \implies (y)^2 -(2)^2\]This is a difference of squares that follows the format \(a^2-b^2 = (a-b)(a+b)\)

OpenStudy (jhannybean):

And youg ot it \(\checkmark\)

OpenStudy (jhannybean):

#3. \(x^2 - 14x - 15\) Find two numbers that add to give you -14, and multiply to give you -15.What would those be?

OpenStudy (anonymous):

#3 is -12-x2-14x right?

OpenStudy (jhannybean):

Not quit. Did you find the factors of -14 and -15?

OpenStudy (jhannybean):

hint: -15 plus what will give you -14?